OpenStudy (anonymous):

Find the slope of the line 2x=7y-11

OpenStudy (anonymous):

II. In each case, find a polynomial equation f(x)=0 satisfying the given conditions. If there is no such equation, say so. 1. the numbers -1, 4, and 5 are roots.

OpenStudy (anonymous):

y=mx+b make your equation look like that form, where y is alone on one side.

OpenStudy (anonymous):

m will be your slope. To illustrate, if I had an equation, 4x=9y-3, I would make it look like y=(4x+3)/9 where my 'm' (slope) would be 4/9

OpenStudy (anonymous):

y= (2x+11)/7

OpenStudy (anonymous):

4. 6x5 +5x4 -29x3 -25x2 -5x =0; roots x=√5, x= -1/3

OpenStudy (anonymous):

Thanks @mabusamha & @monokerous @jenny888 I don't really understand what you were showing me but thank you nonetheless. I am just in basic maths :)